wiki:UC20181029

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 10/29/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Misty Bentz (GSU), Adam Kowalski (CU), Chip Kobulnicky (UWy), Eric Bellm (UW), Ben Williams (UW), Kal Kadlec (UW), Anne Verbiscer (UVa), Sarah Tuttle (UW), Joanne Hughes (Seattle U), Leslie Hebb (NAPG), Russet McMillan (APO), Moire Prescott (NMSU), John Wisniewski (OU)


User feedback and comments from institutional representatives:

  • Seattle - nothing to report
  • NAPG - nothing to report
  • Wyoming - nothing to report
  • Oklahoma - nothing to report
  • Georgia State - nothing to report
  • UVa - one user had a question about the difference between an instrument status being "unchanged" vs. "operational" in the monthly Telescope and Instruments Report - Nancy will clarify this in last month's meeting minutes (DONE)
  • Colorado - nothing to report
  • JHU - no report
  • NMSU - nothing to report
  • Washington - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 9/28/18 through 10/24/18

0) Overview

It has been one of the wettest Octobers in southern NM ever. Observing has been severely limited. Due to a visiting instrument we were able to take DIS down and address a couple of issues; see below for more details.

1) Telescope

The telescope, while not on the sky much, has had no problems.

2) Instruments

DIS: The red camera was installed but required some shims to be reinstated to focus near where the collimator wanted to be for proper alignment. The camera is currently back together and on the vacuum pump. While the dewar was open Bill and John Wilson went through it carefully looking for leaks. The glass interface is still suspect but a couple of scratches were discovered in o-ring grooves. We also had a design review and test fitted a getter chamber that can be ready to install next year.

TripleSpec is operational.

Agile is operational.

Echelle is operational.

NIC-FPS is operational.

ARCTIC is operational.

3) ARCSAT

In use for science.


Additional telescope and instrument discussion:

October has been one of the worst months on record in terms of moisture, and this wetter weather pattern may continue due to El Nino. The main activity for this month has centered on DIS. The engineering staff is taking advantage of a guest instrument being on site to take DIS offline to do some repairs. They are reinstalling some shims that will allow us to focus the CCD with the collimator in a better position; the position has to be changed sometimes due to ambient temperature variations. We are also addressing the contamination. We opened up camera and looked for obvious causes of possible vacuum leaks; nothing major stands out so far. The engineering team did a quick PDR for a getter that would be inserted into the dewar to absorb water vapor and reduce contamination on the flat field lens; this is now in the works. The instrument is running on the vacuum pump now and after final checkout today it will be taken back up to the observing level with the goal of having it back on the sky by Friday.


Q4 3.5-m scheduling:

The Q4 schedule is complete and online. There is some open and DD time available starting in a week. Most of it is B half time, and there is also some time between Christmas and New Years. Users are encouraged to request this time, with the caveat that priority will be given to people who were weathered out in October.


Q4 ARCSAT scheduling:

The Q4 ARCSAT schedule is not completely full yet but it is full until the week of and after Christmas. The Q1 2019 schedule will be done after the Q1 2019 schedule for the 3.5m, but users can request that last open week of Q4 at any time.


Q1 3.5-m scheduling:

Ben will be determining the institutional allocations this week and sending them out to the institutional schedulers late this week. Hopefully by then we will have more news about the status of DIS.


NA2 Baffle Project:

Nancy summarized the current project status. An integrated SolidWorks and Zemax model of the telescope has been completed. The next step is for the UVa team to finalize their initial design of the baffle and then we will contract with an optical design firm to do a detailed analysis of the proposed baffle design.


KOSMOS Update:

The instrument is in the UW lab and the TEG is looking at several different classes of modifications. These include a dewar with a longer hold time, mechanical mods needed to mount the instrument at the NA2 port, a slit viewer camera, and internal calibration lamps. A PDR was held on 10/25/19 and the proposed modifications were presented and discussed in detail. The PDR was successful and the plan is to move forward on a fairly aggressive schedule with the goal of commissioning KOSMOS on the 3.5m in early Q2 2019.


Response to Call for White Papers for 2020 Decadal Survey

Nancy discussed the fact that the current call for Astro2020 white papers (deadline in Jan. 2019) is really focused on science. There will likely be a subsequent call for white papers that deal with facilities, instrumentation, the state of the profession, etc., and that would be a logical call to respond to with a white paper concerning the role of 3-4m class telescopes in the modern era. However, some UC reps felt that we should still prepare something for this white paper call dealing with the ground-based followup of transient sources from TESS, ZTF, LSST, gravitational wave discoveries, etc. Transient followups fall into multiple categories of the science white paper topics, and hence could be neglected because they transcend the topical boundaries. Leslie indicated that she and collaborators are putting together proposals for followup observations of TESS discoveries with ARCTIC + diffuser. Sarah urged Nancy to move forward with a science white paper on this topic and agreed to help.


ARC Board of Governors Meeting

The ARC BoG meeting is on Monday November 5. Nancy will be presenting her Director's Report(s) in a slightly different format this year, with more materials provided in advance for reference (including science highlights, telescope usage, seeing statistics) and allowing for more in-depth discussion of items that the Board can weigh in on (e.g. recommendations from the TDAWG and TEWG). Nancy will debrief this group after the Board meeting.


ACTION ITEMS:

Open action items from previous meetings:

  • Russet will update the calibration section of the ARCTIC users manual once the Q3/Q4 scheduling tasks are completed. STATUS: OPEN. But this is now at the top of her to-do list!
  • UC reps should remind their users who were displaced by DIS unavailability that there is some OPEN time remaining in the Q4 schedule that they can ask for. STATUS: OPEN.
  • Nancy will send out a draft of a white paper for the 2020 Decadal for UC feedback prior to the next telecon. STATUS: OPEN.

New action items from this meeting:

None.


Next meeting:

The next Users Committee telecon will be on Monday December 3 at 10:30 am MST.

Note: See TracWiki for help on using the wiki.